OpenStudy (yrelhan4):

(1 - x^3)^n integration. limits 0 to 1 ??

OpenStudy (experimentx):

try trig sub x^3 = sin^2u

OpenStudy (yrelhan4):

hmm.. i'll try it once and get back to you..

hartnn (hartnn):

if that doesn't work, do u= 1-x^3

OpenStudy (yrelhan4):

how ould that help? @hartnn

hartnn (hartnn):

du = -3x^2 dx du = - 3(1-u)^(2/3)dx du/[- 3(1-u)^(2/3)] = dx just trying.......

OpenStudy (anonymous):

intigrate with respect to x or n?

OpenStudy (experimentx):